### v1

@@ 7.2 Improving Model Explainability
With this work, we propose a method that improves fairness in machine learning models and allows practitioners in various fields to interpret better the decision-making process of the models they utilize. We understand that reducing model bias may not be sufficient in eliminating fairness concerns in decision contexts associated with certain fields such as predictive healthcare, which can be complex, unintuitive, and often hard to explain Yang (2022). In cases like these, solutions must strive for better explainability to improve understanding of the model, increase the utility of its output, and produce better outcomes Yang (2022). This is the reason our work introduces a possible way to strive for more transparency in the process by demonstrating which features affect model prediction the most and how our proposed method influences the model to avoid disparate treatment of unprivileged groups. This also corresponds with the fact that variables or features can and do correlate with protected labels such as race, age, etc. As mentioned in Paulus and Kent (2020), the use of “high dimensional” or “black box” prediction techniques that are typically favored in the machine learning community is generally problematic given that these approaches can predict the protected label through other variables, whether or not the actual protected label is explicitly encoded. This concern is eliminated with our work since our method specifically works by relying less on features connected to the protected label.
In our experiments involving the ADULT dataset, our visualization in Figure 6 showed how our method shifted importance from features that highly correlate with a particular protected label to reduce its effect on the final target label prediction.
For the MIMIC-III experiments, the saliency maps for the baseline model and the model after our proposed method was implemented are also listed in Appendix C. The deeper red color in these figures represents features that hold more weight and, therefore, are more critical for target prediction. Evaluating the map for the baseline model gives us an insight into some of the features that carried the most weight for IHM prediction. For example, the Glascow coma scale (GCS) motor response for Abnormal flexion and extension appears to be highly correlated with IHM prediction. GCS is a tool used to assess a patient’s level of consciousness after a brain injury Jain and Iverson (2018). The GCS motor response score evaluates the patient’s ability to obey commands or move in response to a stimulus. Numerous studies have examined the relationship between the GCS motor response score and in-hospital mortality Yumoto et al. (2019); Sadaka et al. (2012); Settervall et al. (2011). Generally, lower motor response scores are associated with a higher risk of mortality which explains the relevance of these features in our IHM prediction model. Two other features that are shown to be highly correlated with the baseline IHM prediction are Fraction inspired oxygen (FiO2), which is a measure of the concentration of oxygen that a patient is receiving, usually through supplemental oxygen therapy Fuentes and Chowdhury (2020), and pH, which is a measure of the acidity or alkalinity of a patient’s blood Arias-Oliveras (2016). Evidence suggests that higher FiO2 levels may be associated with an increased risk of in-hospital mortality de Jonge et al. (2008), where higher FiO2 levels were associated with an increased risk of mortality in both adult and pediatric patients. There is also some evidence from prior studies that indicate an association between abnormal pH levels and an increased risk of in-hospital mortality Park et al. (2020a), where both low and high pH levels were associated with increased mortality in critically ill patients.
Some features the baseline saliency map shows to carry the least weight in the baseline IHM prediction include diastolic blood pressure, oxygen saturation, and temperature. After implementing our proposed method, the resulting saliency map shows weight shifts and better weight distribution among the features. However, oxygen saturation still carries little weight in the final IHM prediction.
For experiments on SHAPSHOT data, the saliency maps listed in Appendix C give us insights into the features most important for the initial evening-sad-happy target label prediction. The deeper red color in these figures represents features that hold more weight and are, therefore, more important for target prediction across the 7 timesteps. From the saliency maps for the baseline method, a few features show the highest weights in the prediction task, at the top of this list of features being no_sleep_24, phys_3H-10H-medStepsWeightedPeaks, SMS_0H-3H_total_num. Comparing these maps to those from the model produced by our proposed method, it can be noted that our proposed method distributed the weights more evenly among all the features in general, explicitly moving some of the importance away from the more highly-weighted features in the baseline model prediction.
This trend of shifts in feature weights and importance across our experiments agrees with our hypothesis that a shift in feature weight importance to features that correlate less with the protected labels reduces the possibility of the model amplifying or creating biased connections between the protected label and target label.
From our initial work in Zanna et al. (2022), we attributed the notable drop in performance in the previous experiments using our method to the fact that both the target label and chosen protected labels shared similar important features, meaning that shifting weights from these features would have a significant effect on the model’s ability to efficiently form a pattern for the target label, ultimately affecting its prediction. Our results in this work prove this hypothesis to be true. From the visualization of the feature importance seen for ADULT, MIMIC-III, and SNAPSHOT, the models could utilize features that did not highly correlate with the protected label and still maintain an acceptable level of performance for the target label.
@@
The explainable characteristic of our work also eliminates the need for researchers to initially identify the undesired correlation between the target label and the protected label to be able to tackle bias in the model using general existing bias mitigation methods Wang et al. (2020). We hypothesize that our technique can be implemented as a preventative method without knowledge of prior bias in the model. However, this is only applicable to models where our method does not diminish the performance to a detrimental level.
@@ 7.3 Performance-Fairness Trade-off
With the non-dominated sorting part of our method, we can enhance the performance-fairness tradeoff using the concept of Pareto optimality from multi-objective optimization. With this, we are able to obtain a set of Pareto-optimal solutions that give us the ideal models to choose from. With this method, we are providing flexibility for the proprietor to choose what objective, whether fairness or performance, they would like to prioritize and to what extent for every given solution, which can vary according to application or type of data. In our work, we are able to utilize this feature of our method to obtain models with the highest fairness scores since the variability in performance scores was generally low across all experiments.
@@ 7.4 Limitations
One limitation of this research is the restriction to binary outcomes imposed by the tools and techniques developed so far for bias mitigation. The implementation of the existing fairness analysis algorithms provided in the AIF360 toolkit created by IBM requires a binary label dataset as the input, which in models with non-binary labels, might be a barrier to adequately analyzing fairness.
Another limitation is the general lack of an agreed-upon definition of fairness in the field Park et al. (2020b). In the area of Machine Learning, most attempts at proposing new fairness constraints for algorithms have come from the West, and most of the works use the same datasets and problems to show how their constraints perform, which means that there is still no explicit agreement on which constraints are the most appropriate for which problems Mehrabi et al. (2021). Even though fairness is an incredibly desirable quality in society, it can be surprisingly difficult to achieve in practice. With these challenges in mind, there needs to be more standardization regarding fairness definitions and a more rigorous evaluation of what definitions should be applied to different data and problem types. In the future, it will be interesting to determine how popularly-used definitions of fairness translate to models built for data from different domains and how effective they are for each of these models.
Working with deep learning models, we have also encountered limitations associated with the computational complexity of using Monte Carlo dropout. Using this method along with the aspect of saving and loading model weights results in a higher than normal computational cost and time.
@@ 8 Conclusion
Our proposed method introduces a technique for bias mitigation that utilizes feature weights. We demonstrate how this method applies to various kinds of data from 3 different jurisdictions, successfully improving the fairness in the models regardless of the complexity of the predictor’s model. Our method improves fairness in the ADULT, MIMIC-III, and SNAPSHOT datasets at a greater or comparable rate (in the case of SNAPSHOT) to Reweighing, which is a standard bias mitigation technique without sacrificing performance to a substantial level.
Our method also provides an avenue for better model explainability/transparency through saliency maps to elaborate the shifts in feature importance in model prediction tasks. This is a principal characteristic of our work, as transparency is necessary to gain the trust of professionals in various industries, such as health care to have them adopt the predictive models into their daily operations Yang (2022).
Appendix A
In this appendix we provide further description of the ADULT, MIMIC-III and SNAPSHOT datasets from Section 4:
A.1 ADULT Data Description
The ADULT dataset was drawn from the 1994 United States Census Bureau data and involves using personal details such as education level to predict whether an individual will earn more or less than \$50,000 per year Kohavi et al. (1996). This dataset has been extensively utilized in machine learning bias and fairness research Zafar et al. (2017); Friedler et al. (2019); Kearns et al. (2019); Du et al. (2021); Fabris et al. (2022). The extracted features from the dataset used in this work are summarized in table 6.
@@ A.1 ADULT Data Description
A.2 MIMIC-III Data Description
@@
The MIMIC-III (‘Medical Information Mart for Intensive Care’) data set is a large, freely-available database that contains de-identified health-related data associated with over 40,000 patients who stayed in critical care units of the Beth Israel Deaconess Medical Center between 2001 and 2012. MIMIC-III contains detailed information regarding the care of patients, which includes vital signs, medications, laboratory measurements, observations and notes charted by care providers, fluid balance, procedure codes, diagnostic codes, imaging reports, hospital length of stay, survival data, patient demographics and in-hospital mortality (IHM), laboratory test results, billing information, etc. The complete data description can be found in the publication Johnson et al. (2016).
@@
The risk of mortality is often formulated as binary classification using observations recorded from a limited window of time following admission Harutyunyan et al. (2019), and for our experiments, we predicted IHM as a binary label, so the target label indicates whether the patient died before hospital discharge. We processed and cleaned the data according to the method in Harutyunyan et al. (2019), where there are two major steps in the data processing pipeline. First, the root cohort is extracted based on the following exclusion criteria: Hospital admissions with multiple ICU stays or ICU transfers are excluded to reduce any possible ambiguity between outcomes, and ICU stays. Patients younger than 18 are excluded as well. Lastly, event entries are only retained if they can be assigned to a hospital and ICU admission and are part of the list of 17 physiological variables used for modeling, namely: Capillary refill rate, Diastolic blood pressure, Fraction inspired oxygen, Glascow coma scale, eye opening, motor response, verbal response, glucose, heart rate, height, mean arterial pressure, Oxygen saturation, respiratory rate, Systolic blood pressure, temperature, weight, and pH Röösli et al. (2022). From the root cohort, the IHM cohort was extracted by further excluding all ICU stays for which "length of stay" is unknown or less than 48 hours or for which there are no observations in the first 48 hours. The complete process of patient filtering is described in Figure 11 Harutyunyan et al. (2019). This provided final training and test sets of 17,903 and 3,236 ICU stays, respectively. Similar to Harutyunyan et al. (2019), We determined IHM by comparing patient date of death with hospital admission and discharge times. The resulting mortality rate is 13.23% (2,797 of 21,139 ICU stays) Harutyunyan et al. (2019).
@@
The MIMIC-III dataset has some missing data in some instances. For these, we imputed the missing values using the most recent measurement value if it exists and a pre-specified “normal” value otherwise. These pre-specified values are provided by the authors in Harutyunyan et al. (2019). For categorical values, we encoded them using a one-hot vector and standardized the numeric inputs by subtracting the mean and dividing by the standard deviation. We calculated the statistics per variable after the imputation of missing values. After inputting the missing values and standardizing, we obtained 17 pairs of time series for each ICU stay. We have a binary target label for IHM for each stay, determining whether or not a patient dies in the hospital.
@@
From this dataset, the demographic data we utilized in our experiments include age, gender, marital status, and insurance, referred to as the protected labels in this paper. Table 7 shows the distribution of patients, ICU stays, and IHM rates in our data based on the different demographic groups. The age distribution is skewed towards older patients, with over 80% above 50 years. As for ethnicity, two-thirds of the patients are non-Hispanic White. In terms of health insurance, According to Röösli et al. (2022), health insurance type is included as a socioeconomic proxy, and this led to us including it in our experiments. In general, three major insurance types are available in the United States: The public programs Medicare and Medicaid, and private insurance providers. Medicare is a federal social insurance program, and anyone over 65 years or with certain disabilities qualifies for it, whereas Medicaid provides health insurance to very low-income children and their families. Insurance data was therefore mapped to the four distinct categories Medicare, Medicaid, private and other insurance. In this dataset, roughly a third of the patients are covered by private insurance, and more than 50% are enrolled in Medicare. The overall IHM rate in the MIMIC-III dataset is 13.23%, and IHM increases with age. As for gender, female patients appear to have a slightly higher IHM risk, and there is also a large range of variability for ethnic and socioeconomic groups.
@@
A.3 SNAPSHOT Data Description
The SNAPSHOT dataset includes multi-modal physiological, behavioral, and survey data collected from 350 college students in one university, totaling over 7500 days of data Taylor et al. (2017). The study was conducted between 2013 and 2017, within which 30 days of consecutive data were recorded for 228 participants and 90 days for 15 participants. The data recorded includes wrist-worn wearable sensor data (acceleration, skin conductance, and skin temperature), mobile phone data (call, SMS, screen on/off logs, GPS locations), weather data (obtained using DarkSky API Dark Sky (2015)), self-reported daily morning and evening well-being (non-numeric scales or mood, health, and stress later scored 0-100), genders, and Big Five Personality scores John et al. (1999).
From the SNAPSHOT data, we computed a total of 378 features, including 173 physiological features that include Electrodermal activity (EDA) features, skin conductance Level (SCL), skin temperature features, accelerometer features, 165 mobile phone features, 43 weather features from the raw data, and statistical features calculated from each of these classes of features. Tables 8, 9, 10, 11, 12, and 13 summarize these features. For more details on the features extracted, refer to Taylor et al. (2017).
@@ A.3 SNAPSHOT Data Description
State Score
State anxiety score
no_sleep_24
Number of times participant slept in 24 hours.
sleep_latency
Time taken to fall asleep
pre_sleep_activity
What activity the participant performed before going to sleep.
awakening
If the participant woke up through the night (yes or no)
awakening_occations
Number of awakenings
wake_reason
How the participant woke up (spontaneously, alarm, or disturbance).
count_awakening
How many times participant woke up through the night.
awakening_duration
How long was the participant awake for if they woke up through the night.
nap
If the participant took a nap (yes or no)
nap_occations
Number of naps
count_nap
How many times the participant took a nap.
nap_duration
How long the nap was.
first_event_none
If an event is scheduled that day (yes or no)
time_in_bed
How long the participant spent in bed for the day.
sleep_try_time_mins_since_midnight
What time the participant tries to go to sleep
wake_time_mins_since_midnight
Wake time in minutes since midnight
first_event_mins_since_midnight
first event time in minutes since midnight
presleep_media_interaction
If the participant has presleep media interaction (yes or no)
presleep_personal_interaction
If the participant has presleep personal interaction (yes or no)
positive_interaction
If the participant had any positive interactions with someone for the day (yes or no).
negative_interaction
If the participant had any negative interactions with someone for the day (yes or no).
@@
academic
If the participant attended any academic activities (yes or no).
count_academic
How many academic activities the participant attended.
academic_duration
How long the participant spent on academic activities
study_duration
How many hours participant studied for, outside of academic activities.
exercise
If the participants engaged in any exercise-based activities (yes or no)
exercise_occations
How many times the participant engaged in exercise-based activities.
exercise_duration
For how long the participant exercised.
extracurricular
If the participant attended any other extracurricular activities (yes or no).
count_extracurricular
How many extracurricular activities the participant attended in the day.
extracurricular_duration
How long the participant attended extracurricular activities for.
overslept
If the participant overslept and missed any scheduled events.
caffeine_count
Total servings of caffeine participant had for the day.
drugs
If the participant had any other drugs or medication besides (yes or no).
drugs_alcohol
If the participant had any alcohol (yes or no).
drugs_alert
If the participant had any drugs to keep them alert (yes or no).
drugs_sleepy
If the participant had any drugs that made them sleepy (yes or no).
drugs_tired
If the participant had any drugs that made them tired (yes or no).
@@
Sum AUC
Sum of the AUC of all peaks for this period where the amplitude of the peak is calculated as the difference from base tonic signal
Sum AUC Full
Sum of AUC of peaks where amplitude is calculated as difference from 0
Median RiseTime
Median rise time for peaks (seconds)
Median Amplitude
Median amplitude of peaks ($\mu$S)
Count Peaks
Number of detected peaks
SD Peaks 30 min
Compute number of peaks per 30 minute epoch, take standard deviation of this signal
Med Peaks 30 min
Compute number of peaks per 30 minute epoch, take median of this signal
Percent Med Peak
Percentage of signal containing 1 minute epochs with greater than 5 peaks
Percent High Peak
Same as Percent Med Peak
@@
Percent Off
Percentage of period where sensor was off
Max Unnorm
Maximum level of un-normalized EDA signal
Med Unnorm
Median of un-normalized EDA signal
Mean Unnorm
Mean of un-normalized EDA signal
Median Norm
Median of z-score normalized EDA signal
SD Norm
Standard Deviation of z-score normalized EDA signal
Mean Deriv
Mean derivative of z-score normalized EDA signal ($\mu$S/second)
@@
Step Count
Number of steps detected
Mean Movement Step Time
Average number of samples (at 8Hz) between two steps (aggregated first to 1 minute, then we take the mean of only the parts of this signal occurring during movement)
Stillness Percent
Percentage of time the person spent nearly motionless
Sum Stillness weighted AUC
Sum the weights of the peak AUC signal by how still the user was every 5 minutes
Sum Steps Weighted AUC
Sum the weights of the peak AUC signal by the step count over every 5 minutes
Sum Stillness Weighted Peaks
Multiply the number of peaks every 5 minutes by the amount of stillness during that period
Max Stillness Weighted Peaks
Max value for the number of peaks multiplied by the stillness for any five minute period
Sum Steps Weighted Peaks
Divide number of peaks every five minutes by step count and sum
Med Steps Weighted Peaks
Average value of number of peaks divided by step count for every 5 mins
@@
Max Raw Temp
Maximum of the raw temperature signal (°C)
Min Raw Temp
Minimum of the raw temperature signal (°C)
SD Raw Temp
Standard deviation of the raw temperature signal
Med Raw Temp
Standard deviation of the raw temperature signal
Sum Temp Weighted AUC
Sum of peak AUC divided by the average temp for every 5 minutes
Sum Temp Weighted Peaks
Number of peaks divided by the average temp for every 5 minutes
Max Temp Weighted Peaks
Maximum number of peaks in any 5 minute period divided by the average temperature
SD Stillness Temp
Standard deviation of the temperature recorded during periods when the person was still
Med Stillness Temp
Median of the temperature recorded during periods when the person was still

### v2

@@ 7.2 Improving Model Explainability
In this work, we propose a method that not only enhances fairness in machine learning models but also aids practitioners across various sectors in better understanding the decision-making processes of their models. Recognizing that simply reducing model bias may not fully address fairness concerns, especially in complex fields like predictive healthcare which can be complex, unintuitive, and often hard to explain, our method focuses on improving model explainability to increase the utility of outputs and achieve better outcomes (Yang (2022)). By demonstrating which features most influence model predictions and adjusting feature weight away from protected labels, our approach promotes transparency and mitigates disparate treatment of unprivileged groups.
Our experiments demonstrate these concepts effectively. For instance, using the ADULT dataset, Figure 4 indicates our method reduces reliance on features closely associated with protected labels, thereby minimizing their impact on predictions. Similarly, saliency maps for the MIMIC-III and SNAPSHOT datasets, detailed in Appendix C, show our method’s ability to redistribute feature importance, ensuring a more balanced consideration of inputs unrelated to protected attributes. This alignment with our hypothesis confirms that de-emphasizing features correlated with protected labels diminishes bias without significantly compromising model performance.
@@
Additionally, the explainable characteristic of our work also eliminates the need for researchers to initially identify the undesired correlation between the target label and the protected label to be able to tackle bias in the model using general existing bias mitigation methods (Wang et al. (2020)).
@@ 7.3 Performance-Fairness Trade-off
Our method employs non-dominated sorting and the concept of Pareto optimality from multi-objective optimization to enhance the performance-fairness trade-off. This approach has allowed us to identify a set of Pareto-optimal solutions, presenting a range of models from which proprietors can select based on their specific needs. This method offers flexibility, enabling users to prioritize either fairness or performance according to the application or data type. In our experiments, we leverage this feature to select models with the highest fairness scores, as the variability in performance is consistently low across all tested scenarios.
@@ 7.4 Limitations
One limitation of this research is the restriction to binary outcomes imposed by the tools and techniques utilized for bias mitigation and the outcome labels. Another limitation is the general lack of an agreed-upon definition of fairness in the field (Park et al. (2020b)). In the area of Machine Learning, most attempts at proposing new fairness constraints for algorithms have come from the West, and most of the works use the same datasets and problems to show how their constraints perform, which means that there is still no explicit agreement on which constraints are the most appropriate for which problems (Mehrabi et al. (2021)). Even though fairness is an incredibly desirable quality in society, it can be surprisingly difficult to achieve in practice. With these challenges in mind, there needs to be more standardization regarding fairness definitions and a more rigorous evaluation of what definitions should be applied to different data and problem types. In the future, it will be interesting to determine how popularly-used definitions of fairness translate to models built for data from different domains and how effective they are for each of these models in real-world scenarios.
Working with deep learning models, we have also encountered limitations associated with the computational complexity of using MC Dropout. Using this method along with the aspect of saving and loading model weights results in a higher than normal computational cost and time.
@@ 8 Conclusion
Our proposed method leverages feature weights for bias mitigation and has been successfully applied to data from three different jurisdictions—enhancing fairness across the ADULT, MIMIC-III, and SNAPSHOT datasets. It achieves improvements in fairness at a rate comparable to or greater than that of Reweighing, a standard bias mitigation technique, and other existing methods, without significant sacrifices in performance. Additionally, our approach offers flexibility in managing the fairness-performance trade-off, utilizing principles from multi-objective optimization to allow users to prioritize their desired objectives.
Furthermore, our method enhances model transparency through the use of saliency maps, which detail shifts in feature importance for prediction tasks. This transparency is crucial for building trust among professionals in various fields, such as healthcare, encouraging them to integrate these predictive models into their daily operations (Yang (2022)).
Appendix A. Data Description
In this appendix we provide further description of the ADULT, MIMIC-III and SNAPSHOT datasets and our data processing procedures.
.1 ADULT Data
The ADULT dataset was drawn from the 1994 United States Census Bureau data and involves using personal details such as education level to predict whether an individual will earn more or less than \$50,000 per year (Kohavi et al. [1996]). This dataset has been extensively utilized in machine learning bias and fairness research (Zafar et al. [2017], Friedler et al. [2019], Kearns et al. [2019], Du et al. [2021], Fabris et al. [2022]). The extracted features from the dataset used in this work are summarized in Table 6.
@@ Appendix A. Data Description
This dataset contains 14 categorical and continuous integer attributes from which we dropped columns with noisy data (over 100 missing values). We were left with 8 features: Age, Workclass, Education, Marital status, Occupation, Relationship, Race, and Sex.
The prediction task for this dataset is a binary income/salary prediction, where a label of 1 represents participants who earn $>$ 50k and 0 for those who earn <=<=< = 50k in a year. We converted all the string features to integers, and we used 45222 of 48842 instances after dropping a total of 3620 samples with noisy data. We utilized age, sex (gender), and race as the protected labels in our experiments.
For fairness analysis, we binarized the 3 protected labels to 0 as unprivileged and 1 as privileged classes. For the age label, 12% of the female class earned more than 50k in salary, while for the male class, 32% earned more than 50k. 36% of participants above the age of 40 earned more than 50k, while 17% of those below the age of 40 earned more than 50k. As for the race label, less than 13% of the participants who identified as Black, Amer-Indian-Eskimo, and "Other" earned more than 50k, while about 30% of those who identified as White or Asian-Pacific -Islander earned more than 50k. We chose the class corresponding to higher income/salary as the privileged class for each protected label. These privileged classes are also the majority classes in the dataset. For the age label, since the difference in the number of participants between classes is larger than the difference in average salary earned, we chose the class with more participants (less than or equal to 40) as the privileged class (1), while the class with fewer participants (over 40) as the unprivileged class (0). For race, Black(0), Amer-Indian-Eskimo(4), and other(2) became 0, and White(3) and Asian-Pac-Islander(1) became 1, and for sex, male became 1 and female became 0. We excluded the protected labels as features to feed into our prediction model, and split the data into 75% training and 25% testing sets.
.2 MIMIC-III Data Description
@@
The MIMIC-III (‘Medical Information Mart for Intensive Care’) data set is a large, freely-available database that contains de-identified health-related data associated with over 40,000 patients who stayed in critical care units of the Beth Israel Deaconess Medical Center. It contains detailed information regarding the care of patients, which includes vital signs, medications, laboratory measurements, observations and notes charted by care providers, fluid balance, procedure codes, diagnostic codes, imaging reports, hospital length of stay, survival data, patient demographics and in-hospital mortality (IHM), laboratory test results, billing information, etc. The complete data description can be found in the publication (Johnson et al. [2016]).
@@
The risk of mortality is often formulated as binary classification using observations recorded from a limited window of time following admission (Harutyunyan et al. [2019]), and for our experiments on MIMIC-III, we predicted IHM as a binary label, so the target label indicates whether the patient died before hospital discharge. We processed and cleaned the data according to the method in Harutyunyan et al. [2019], where there are two major steps in the data processing pipeline. First, the root cohort is extracted based on the following exclusion criteria: Hospital admissions with multiple ICU stays or ICU transfers are excluded to reduce any possible ambiguity between outcomes, and ICU stays. Patients younger than 18 are excluded as well. Lastly, event entries are only retained if they can be assigned to a hospital and ICU admission and are part of the list of 17 physiological variables used for modeling, namely: Capillary refill rate, Diastolic blood pressure, Fraction inspired oxygen, Glascow coma scale, eye opening, motor response, verbal response, glucose, heart rate, height, mean arterial pressure, Oxygen saturation, respiratory rate, Systolic blood pressure, temperature, weight, and pH (Röösli et al. [2022]). From the root cohort, the IHM cohort was extracted by further excluding all ICU stays for which "length of stay" is unknown or less than 48 hours or for which there are no observations in the first 48 hours. The complete process of patient filtering is described in Figure 9 (Harutyunyan et al. [2019]). This provided final training and test sets of 17,903 and 3,236 ICU stays, respectively. Similar to Harutyunyan et al. [2019], We determined IHM by comparing patient date of death with hospital admission and discharge times. The resulting mortality rate is 13.23% (2,797 of 21,139 ICU stays) (Harutyunyan et al. [2019]).
@@
The MIMIC-III dataset has some missing data in some instances. For these, we imputed the missing values using the most recent measurement value if it exists and a pre-specified “normal” value otherwise. These pre-specified values are provided by the authors in Harutyunyan et al. [2019]. For categorical values, we encoded them using a one-hot vector and standardized the numeric inputs by subtracting the mean and dividing by the standard deviation. We calculated the statistics per variable after the imputation of missing values. After inputting the missing values and standardizing, we obtained 17 pairs of time series for each ICU stay. We have a binary target label for IHM for each stay, determining whether or not a patient dies in the hospital.
@@
From this dataset, the demographic data we utilized in our experiments include age, gender, marital status, and insurance, referred to as the protected labels in this paper. Table 7 shows the distribution of patients, ICU stays, and IHM rates in our data based on the different demographic groups. The age distribution is skewed towards older patients, with over 80% above 50 years. As for ethnicity, two-thirds of the patients are non-Hispanic White. In terms of health insurance, According to Röösli et al. [2022], health insurance type is included as a socioeconomic proxy, and this led to us including it in our experiments. In general, three major insurance types are available in the United States: The public programs Medicare and Medicaid, and private insurance providers. Medicare is a federal social insurance program, and anyone over 65 years or with certain disabilities qualifies for it, whereas Medicaid provides health insurance to very low-income children and their families. Insurance data was therefore mapped to the four distinct categories Medicare, Medicaid, private, and other insurance. In this dataset, roughly a third of the patients are covered by private insurance, and more than 50% are enrolled in Medicare. The overall IHM rate in the MIMIC-III dataset is 13.23%, and IHM increases with age. As for gender, female patients appear to have a slightly higher IHM risk, and there is also a large range of variability for ethnic and socioeconomic groups.
@@
With this dataset, we predicted in-hospital mortality (IHM) as a binary label. We determined IHM by comparing the patient’s date of death with hospital admission and discharge times. The resulting mortality rate is 13.23% (2,797 of 21,139 ICU stays). We processed data by reproducing the method in Harutyunyan et al. [2019]. A full description can be found in the original publication. To process the data for fairness evaluation, we binary-encoded each protected label and assigned them as privileged and unprivileged classes to analyze them using our chosen fairness metrics. We encoded all participants with age $<=60$ as the unprivileged class (0), and those $>60$ as the privileged class (1). The reasoning behind this encoding is that the Centers for Disease Control and Prevention (CDC) defines an “older adult” as someone who is at least 60 years old (for Disease Control et al. [2012]). For the Gender label, male was encoded as the privileged class (1) due to its higher number of samples and because female patients appear to have a slightly higher IHM risk, making the female class the unprivileged class (0). We classified the participants according to whether they have private or government-assisted insurance. Those with private or self-care insurance represented the privileged class (1), and those with Medicare, Medicaid, and Government represented the unprivileged class (0). The reasoning behind this is the connection between government-aided insurance and socioeconomic conditions and age as mentioned earlier in this section. For marital status, we categorized married patients, those with life partners as privileged, and all others (no partner) as unprivileged. We did this in accordance with research showing that unmarried patients, including those who are divorced, separated, widowed, or never married, show elevated rates of certain health complications and death compared to those with partners (Dhindsa et al. [2020], Johnson et al. [2000]).
.3 SNAPSHOT Data Description
The SNAPSHOT dataset includes multi-modal physiological, behavioral, and survey data collected from 350 college students in one university, totaling over 7500 days of data (Taylor et al. [2017]). The study was conducted between 2013 and 2017, within which 30 days of consecutive data were recorded for 228 participants and 90 days for 15 participants. The data recorded includes wrist-worn wearable sensor data (acceleration, skin conductance, and skin temperature), mobile phone data (call, SMS, screen on/off logs, GPS locations), weather data (obtained using DarkSky API Dark Sky [2015]), self-reported daily morning and evening well-being (non-numeric scales or mood, health, and stress later scored 0-100), genders, and Big Five Personality scores (John et al. [1999]).
From the SNAPSHOT data, we computed a total of 378 features, including 173 physiological features that include Electrodermal activity (EDA) features, skin conductance Level (SCL), skin temperature features, accelerometer features, 165 mobile phone features, 43 weather features from the raw data, and statistical features calculated from each of these classes of features. Tables 8, 9, 10, 11, 12, and 13 summarize these features. For more details on the features extracted, refer to Taylor et al. [2017].
@@
State Score State anxiety score
no_sleep_24 Number of times participant slept in 24 hours.
sleep_latency Time taken to fall asleep
pre_sleep_activity What activity the participant performed before going to sleep.
awakening If the participant woke up through the night (yes or no)
awakening_occations Number of awakenings
wake_reason How the participant woke up (spontaneously, alarm, or disturbance).
count_awakening How many times participant woke up through the night.
awakening_duration How long was the participant awake for if they woke up through the night.
nap If the participant took a nap (yes or no)
nap_occations Number of naps
count_nap How many times the participant took a nap.
nap_duration How long the nap was.
first_event_none If an event is scheduled that day (yes or no)
time_in_bed How long the participant spent in bed for the day.
sleep_try_time_mins_since_midnight What time the participant tries to go to sleep
wake_time_mins_since_midnight Wake time in minutes since midnight
first_event_mins_since_midnight first event time in minutes since midnight
presleep_media_interaction If the participant has presleep media interaction (yes or no)
presleep_personal_interaction If the participant has presleep personal interaction (yes or no)
positive_interaction If the participant had any positive interactions with someone for the day (yes or no).
negative_interaction If the participant had any negative interactions with someone for the day (yes or no).
@@
academic If the participant attended any academic activities (yes or no).
count_academic How many academic activities the participant attended.
academic_duration How long the participant spent on academic activities
study_duration How many hours participant studied for, outside of academic activities.
exercise If the participants engaged in any exercise-based activities (yes or no)
exercise_occations How many times the participant engaged in exercise-based activities.
exercise_duration For how long the participant exercised.
extracurricular If the participant attended any other extracurricular activities (yes or no).
count_extracurricular How many extracurricular activities the participant attended in the day.
extracurricular_duration How long the participant attended extracurricular activities for.
overslept If the participant overslept and missed any scheduled events.
caffeine_count Total servings of caffeine participant had for the day.
drugs If the participant had any other drugs or medication besides (yes or no).
drugs_alcohol If the participant had any alcohol (yes or no).
drugs_alert If the participant had any drugs to keep them alert (yes or no).
drugs_sleepy If the participant had any drugs that made them sleepy (yes or no).
drugs_tired If the participant had any drugs that made them tired (yes or no).
@@
Sum AUC Sum of the AUC of all peaks for this period where the amplitude of the peak is calculated as the difference from base tonic signal
Sum AUC Full Sum of AUC of peaks where amplitude is calculated as difference from 0
Median RiseTime Median rise time for peaks (seconds)
Median Amplitude Median amplitude of peaks ($\mu$S)
Count Peaks Number of detected peaks
SD Peaks 30 min Compute number of peaks per 30 minute epoch, take standard deviation of this signal
Med Peaks 30 min Compute number of peaks per 30 minute epoch, take median of this signal
Percent Med Peak Percentage of signal containing 1 minute epochs with greater than 5 peaks
Percent High Peak Same as Percent Med Peak
@@
Percent Off Percentage of period where sensor was off
Max Unnorm Maximum level of un-normalized EDA signal
Med Unnorm Median of un-normalized EDA signal
Mean Unnorm Mean of un-normalized EDA signal
Median Norm Median of z-score normalized EDA signal
SD Norm Standard Deviation of z-score normalized EDA signal
Mean Deriv Mean derivative of z-score normalized EDA signal ($\mu$S/second)
@@
Step Count Number of steps detected
Mean Movement Step Time Average number of samples (at 8Hz) between two steps (aggregated first to 1 minute, then we take the mean of only the parts of this signal occurring during movement)
Stillness Percent Percentage of time the person spent nearly motionless
Sum Stillness weighted AUC Sum the weights of the peak AUC signal by how still the user was every 5 minutes
Sum Steps Weighted AUC Sum the weights of the peak AUC signal by the step count over every 5 minutes
Sum Stillness Weighted Peaks Multiply the number of peaks every 5 minutes by the amount of stillness during that period
Max Stillness Weighted Peaks Max value for the number of peaks multiplied by the stillness for any five minute period
Sum Steps Weighted Peaks Divide number of peaks every five minutes by step count and sum
Med Steps Weighted Peaks Average value of number of peaks divided by step count for every 5 mins
@@
Max Raw Temp Maximum of the raw temperature signal (°C)
Min Raw Temp Minimum of the raw temperature signal (°C)
SD Raw Temp Standard deviation of the raw temperature signal
Med Raw Temp Standard deviation of the raw temperature signal
Sum Temp Weighted AUC Sum of peak AUC divided by the average temp for every 5 minutes
Sum Temp Weighted Peaks Number of peaks divided by the average temp for every 5 minutes
Max Temp Weighted Peaks Maximum number of peaks in any 5 minute period divided by the average temperature
SD Stillness Temp Standard deviation of the temperature recorded during periods when the person was still
Med Stillness Temp Median of the temperature recorded during periods when the person was still